# Finite groups and arc-transitive maps of square-free Euler characteristic

### v2

@@ 1. Introduction
As usual, for a group $G$ and a subset $\pi$ of prime divisors of $|G|$, denote by $G_{\pi}$ a Hall $\pi$-subgroup of $G$, which is a Sylow subgroup if $\pi$ contains a single prime $p$, and denoted by $G_{p}$. For a positive integer $n$, denote by ${\sf Z}_{n}$ a cyclic group of order $n$, by ${\rm D}_{2n}$ or ${\rm Q}_{2n}$ a dihedral or generalized quaternion group of order $2n$, and by ${\rm A}_{n}$ or ${\rm S}_{n}$ an alternating or symmetric group of degree $n$, respectively. A finite group $G$ is a semidirect product of $A$ by $B$ if $A\lhd G$ and $B<G$ are such that $G/A\cong B$, denoted by G=A:BG=A{:}B. Our first main result is the following theorem.
@@
(1) The Sylow subgroups of $G$ are characterized in Proposition 4.1 (for $G_{2}$) and Lemma 4.3 (for $G_{p}$, $p$ odd prime).
(2) G=(A:B):KG=(A{:}B){:}K, where $\gcd(|A{:}B|,|K|)=1$, such that
@@
    $K$ is one of the following groups:
    • (a)
      $K=G_{2}$ ($K$ might be $1$ in case of $G$ of odd order);
    • (b)
      $K=G_{\{2,3\}}$, explicitly given in Lemma 5.6 (2) - (3);
    • (c)
      $K=G_{\{2,7\}}={\sf Z}_{2}^{3}{:}G_{7}$, or $K=G_{\{2,3,7\}}={\sf Z}_{2}^{3}{:}(G_{7}{:}G_{3})$.
@@
Let $\mathcal{M}$} be a map with square-free Euler characteristic, and let $G\leqslant{\sf Aut}{\mathcal{M}}$ be solvable and arc-transitive on $\mathcal{M}$}. Then G=(A:B):KG=(A{:}B){:}K such that $A$ is abelian with $A\cap{\bf Z}(A{:}B)=1$, $B$ is nilpotent, and either $K=G_{2}$ is ${\sf Z}_{2}$ or one of the groups given in Proposition 4.2, or one of the following holds:
@@
    $\mathcal{M}$} is $G$-regular, $K=G_{\{2,3\}}$ is one of $X$, $X\times{\sf Z}_{2}$, ${\rm D}_{2.3^{\ell}}\times{\rm S}_{4}$ or $({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})\times{\rm D}_{6}$, where $X\in\{{\rm S}_{4},\,{\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell+1}},\,({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2},\,({\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}\};$
@@
    $\mathcal{M}$} is $G$-vertex-reversing, $K=G_{\{2,3\}}$ is one of the groups in Table 1 and Table 2, corresponding to cases (1.1), (1.2), (1.5), (1.6), (2.4), (2.5);
@@
    $\mathcal{M}$} is $G$-vertex-rotary, $K=G_{\{2,7\}}={\sf Z}_{2}^{3}{:}{\sf Z}_{7^{\ell}}$, or $K=G_{\{2,3\}}$ is one of ${\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}}$, ${\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})$, ${\sf Z}_{4}\circ({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})$, or one of the groups in Table 1 and Table 2, excluding cases (1.5), (1.6).
The proofs of the above two results are given in Section 4 - 5 and Section 6. Before that, we construct several interesting examples of regular maps with square-free Euler characteristic in Section 3, from which the following result follows.
@@
There are infinite families of regular maps $\mathcal{M}$} with underlying graphs ${\bf C}_{n}^{(n)}$ or ${\bf C}_{n}\square{\bf C}_{n}$, and with Euler characteristic $\chi({\mathcal{M}})=-n(n-3)$ or n(n2)-n(n-2), respectively. In particular, there are infinitely many of them with $\chi({\mathcal{M}})$ square-free.
@@ 2. Preliminaries
Let $G$ be a group. Denote by $\Phi(G)$ and $G^{\prime}$ the Frattini subgroup and the commutator subgroup of $G$. If $G$ is a $p$-group, then $^{\prime}\leqslant\Phi(G), G/G$$G^{\prime}$ is abelian, and $G/\Phi(G)$ is elementary abelian. For a subgroup $H\leqslant G$, denote by ${\bf N}_{G}(H)$ and ${\bf C}_{G}(H)$ the normalizer and the centralizer of $H$ in $G$, respectively.
@@
A flag of $\mathcal{M}$} is an incidence triple $(\alpha,e,f)$ with $\alpha\in V$, $e\in E$ and $f\in F$. Let $G\leqslant{\sf Aut}({\mathcal{M}})$, and let $(\alpha,e,f)$ be a flag. It is well-known that each of the stabilizers $G_{\alpha}$ and $G_{f}$ is cyclic or dihedral, and meanwhile, $G_{e}=1$, ${\sf Z}_{2}$ or ${\rm D}_{4}$.
@@
(1) $\mathcal{M}$} is a $G$-regular map (type 1), which is determined a regular triple $(x,y,z)$ for $G$ satisfying:
$\mbox{$G=\langle x,y,z\rangle$, $|x|=|y|=|z|=2$ and $\langle x,z\rangle\cong{\sf Z}_{2}^{2}$}.$
In this case, we have:
$G={\sf Aut}({\mathcal{M}}),\,G_{\alpha}=\langle x,y\rangle,\,G_{e}=\langle x,z\rangle,\,G_{f}=\langle y,z\rangle,$
and the underlying graph of $\mathcal{M}$} can be expressed as a coset graph ${\sf Cos}(G,H,HzH)$, where $H=\langle x,y\rangle$ (refer to [12].)
@@
The following lemma established in [11] characterizes Sylow subgroups of ${\sf Aut}({\mathcal{M}})$ in the case where $\mathcal{M}$} is with square-free characteristic. We note that if $\chi({\mathcal{M}})$ is square-free, then of course so is $\gcd(\chi({\mathcal{M}}),|{\sf Aut}{\mathcal{M}}|)$.
@@ 3. Examples of square-free Euler characteristic regular maps
We construct several infinite families of regular maps with square-free Euler characteristic in this section. The constructions are based on the following groups:
@@
The first family is given now.
@@
Let $G=\langle a,s\rangle\times\langle b,t\rangle={\rm D}_{2n}\times{\rm D}_{2n}\leqslant X$ as given in ($\ast$), with $n$ odd. Let
@@
Then $(x,y,z)$ is a triple of involutions with $xz=zx$. Further,
$(yx)^{2}=(abt)^{2}=a^{2}$ and $(yxz)^{2}=(abs)^{2}=b^{2}$.
As $n$ is odd, $\langle a\rangle=\langle a^{2}\rangle$ and $\langle b\rangle=\langle b^{2}\rangle$, so that $G=\langle x,y,z\rangle$. Thus, $(x,y,z)$ is a regular triple for $G$.
@@
$H=\langle x,y\rangle=\langle s,abst\rangle=\langle abt\rangle{:}\langle s\rangle=\langle a,s\rangle\times\langle bt\rangle={\rm D}_{4n}.$
Meanwhile, the face stabilizer is
@@
Thus the map ${\mathcal{M}}_{n}$ has valency $2n$ and face length $n$. Further, $H\cap H^{z}=\langle a,s\rangle\cong{\rm D}_{2n}$ is of index $2$ in $H$, so the underlying graph ${\it\Gamma}={\sf Cos}(G,H,HzH)\cong{\bf C}_{n}^{(n)}$.
The Euler characteristic
$\chi({\mathcal{M}}_{n})=|G|\left({1\over|\langle x,y\rangle|}-{1\over 4}+{1\over|\langle y,z\rangle|}\right)=4n^{2}\left({1\over 4n}-{1\over 4}+{1\over 2n}\right)=-n(n-3).$
Let $n=4x+1$. Then $\chi({\mathcal{M}}_{n})=-2(4x+1)(2x-1)$. By a result given by G. Ricci in 1933 (which is improved by $\rm P.Erd\ddot{o}s$} in [8]), there are infinitely many positive integers $x$ for which f(x)=(4x+1)(2x1)f(x)=(4x+1)(2x-1) is square-free, so there are infinitely many odd values of $n$ such that $\chi({\mathcal{M}}_{n})$ is square-free. $\Box$
Here is some data such that n(n3)-n(n-3) is square-free, with $n$ odd.
$\begin{array}[]{c|ccccccc}\hline\cr n&5&13&17&29&37&41&\cdots\\ -n(n-3)&-2.5&-2.5.13&-2.7.17&-2.13.29&-2.17.37&-2.19.41$$\cdots\\ \hline\cr\end{array}$
Next we construct the second family.
@@
Let $G=\langle a,b\rangle{:}\langle st,\sigma\rangle={\sf Z}_{n}^{2}{:}{\sf Z}_{2}^{2}\leqslant X$ as given in ($\ast$). Let
@@
Then $(x,y,z)$ is a triple of involutions with $xz=zx$. Further,
$yz=b^{-1}$ and $(yz)^{x}=a^{-1}$,
so that $G=\langle x,y,z\rangle$. Thus, $(x,y,z)$ is a regular triple for $G$.
@@
$H=\langle x,y\rangle=\langle ast\sigma\rangle{:}\langle\sigma\rangle={\rm D}_{4n}.$
Meanwhile, the face stabilizer is
@@
Thus the map ${\mathcal{M}}_{n}$ has valency $2n$ and face length $n$. Further, $H\cap H^{z}=\langle a^{-1}b\rangle{:}\langle\sigma\rangle={\rm D}_{2n}$ is of index $2$ in $H$, so the underlying graph ${\it\Gamma}={\sf Cos}(G,H,HzH)\cong{\bf C}_{n}^{(n)}$.
The Euler characteristic
$\chi({\mathcal{M}})=|G|\left({1\over|\langle x,y\rangle|}-{1\over 4}+{1\over|\langle y,z\rangle|}\right)=4n^{2}\left({1\over 4n}-{1\over 4}+{1\over 2n}\right)=-n(n-3).$
By Lemma 3.2, there are infinitely many odd values of $n$ such that $\chi({\mathcal{M}}_{n})$ is square-free. Further, let $n=4x+2$. Then according to [8], there are infinitely many positive integers $x$ for which f(x)=(2x+1)(4x1)f(x)=(2x+1)(4x-1) is square-free, so there are infinitely many even values of $n$ such that $\chi({\mathcal{M}}_{n})=-2(2x+1)(4x-1)$ is square-free. $\Box$
Here is some data such that n(n3)-n(n-3) is square-free, with $n$ even.
$\begin{array}[]{c|ccccccc}\hline\cr n&2&10&14&22&26&34&\cdots\\ -n(n-3)&2&-2.5.7&-2.7.11&-2.11.19&-2.13.23&-2.17.31$$\cdots\\ \hline\cr\end{array}$
@@
Let $G=\langle a,b\rangle{:}\langle s,\sigma\rangle={\sf Z}_{n}^{2}{:}{\rm D}_{8}=X$ as given in ($\ast$), with $n$ odd. Let
@@
Then $(x,y,z)$ is a triple of involutions with $xz=zx$. Further,
$(zy)^{2}=(abt)^{2}=a^{2}$ and $((zy)^{2})^{x}=(a^{2})^{x}=b^{2}$.
As $n$ is odd, $\langle a\rangle=\langle a^{2}\rangle$ and $\langle b\rangle=\langle b^{2}\rangle$, so that $G=\langle x,y,z\rangle$. Thus, $(x,y,z)$ is a regular triple for $G$.
@@
$H=\langle x,y\rangle=\langle\sigma,s\rangle={\rm D}_{8}.$
Meanwhile, the face stabilizer

### v3

@@ 1. Introduction
As usual, for a group $G$ and a subset $\pi$ of prime divisors of $|G|$, denote by $G_{\pi}$ a Hall $\pi$-subgroup of $G$, which is a Sylow subgroup if $\pi$ contains a single prime $p$, and denoted by $G_{p}$. For a positive integer $n$, denote by ${\sf Z}_{n}$ the cyclic group of order $n$, by ${\rm D}_{2n}$ or ${\rm Q}_{2n}$ the dihedral or generalized quaternion group of order $2n$, and by ${\rm A}_{n}$ or ${\rm S}_{n}$ the alternating or symmetric group of degree $n$, respectively. A finite group $G$ is a semidirect product of $A$ by $B$ if $A\lhd G$ and $B<G$ are such that $G/A\cong B$, denoted by G=A:BG=A{:}B. Denote by $A\circ_{C}B$ the central product of $A$ and $B$ (find the definition in Preliminaries). Our first main result is the following theorem.
@@
(1) The Sylow subgroups $G_{2}$ and $G_{p}\ (p>2)$ of $G$ are respectively given in Proposition 4.1 and Lemma 4.3.
(2) Let $H\leqslant G$ be the largest normal Hall subgroup of odd order. Then, H=A:BH=A{:}B, where $A$ is abelian, $B$ is nilpotent and $A\cap{\bf Z}(H)=1$.
(3) G=H:KG=H{:}K, where $K$ satisfies one of the following :
@@
    $K=G_{\{2,3\}}$ or $G_{\{2,7\}}$ is one of the groups :
    ${\sf Z}_{2}^{2}{:}G_{3},\ {\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}G_{3}),\ {\rm Q}_{8}{:}G_{3},\ {\sf Z}_{4}\circ_{{\sf Z}_{2}}({\rm Q}_{8}{:}G_{3}),\,{\sf Z}_{2}^{3}{:}G_{7}\,;$
  • (iii)
    $K=G_{\{2,3\}}$ is one of the groups :
    $X$, ${\sf Z}_{2}\times X$, $Y$, ${\sf Z}_{4}\circ_{{\sf Z}_{2}}Y$, $W$,
    where $X,Y,W$ are given in Table 1, $X$ is homomorphic to ${\sf Z}_{2}^{2}{:}{\rm S}_{3}\cong{\rm S}_{4}$, $Y$ is homomorphic to ${\rm Q}_{8}{:}{\rm S}_{3}\cong{\rm GL}_{2}(3)$, and $W$ is homomorphic to ${\sf Z}_{2}^{2}{:}{\sf Z}_{3}\cong{\rm A}_{4}$, ${\rm S}_{4}$, ${\sf Z}_{4}\circ{\rm SL}_{2}(3)$ or ${\sf Z}_{4}\circ{\rm GL}_{2}(3);$
    Table 1. $X,Y,W$, $\ell\geqslant 1$
    Row $X$ $Y$ $W$
    11 ${\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}}$ ${\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}}$ ${\rm D}_{2.3^{\ell}}\times{\rm A}_{4}$, $({\sf Z}_{3^{\ell}}{:}{\sf Z}_{4})\circ({\rm Q}_{8}{:}{\sf Z}_{3})$
    22 ${\sf Z}_{3^{\ell}}\times{\rm S}_{4}$ ${\sf Z}_{3^{\ell}}\times{\rm GL}_{2}(3)$ $({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})\times{\rm D}_{6}$, $({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})\circ({\sf Z}_{3}{:}{\sf Z}_{4})$
    33 $({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})\times{\sf Z}_{3}$ $({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})\times{\sf Z}_{3}$ ${\rm D}_{2.3^{\ell}}\times{\rm S}_{4}$, $({\sf Z}_{3^{\ell}}{:}{\sf Z}_{4})\circ({\rm Q}_{8}{:}{\rm S}_{3})$
    44 $({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2}$ $({\sf Z}_{3^{\ell}}\times({\rm Q}_{8}{:}{\sf Z}_{3})){:}{\sf Z}_{2}$ $({\sf Z}_{2}^{2}{:}{{\rm D}_{2.3^{\ell}}})\times{\rm D}_{6}$, $({\rm Q}_{8}{:}{{\rm D}_{2.3^{\ell}}})\circ({\sf Z}_{3}{:}{\sf Z}_{4})$
    55 $(({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})\times{\sf Z}_{3}){:}{\sf Z}_{2}$ $(({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})\times{\sf Z}_{3}){:}{\sf Z}_{2}$ $({\sf Z}_{2}^{2}\times{\rm D}_{2.3^{\ell+1}}){:}{\sf Z}_{3}$, $({\rm Q}_{8}\circ({\sf Z}_{3^{\ell+1}}{:}{\sf Z}_{4})){:}{\sf Z}_{3}$
    66 ${\sf Z}_{2}^{2}{:}({\sf Z}_{3^{\ell+1}}{:}{\sf Z}_{6})$ ${\rm Q}_{8}{:}({\sf Z}_{3^{\ell+1}}{:}{\sf Z}_{6})$
  • (iv)
    $K=G_{\{2,3,7\}}={\sf Z}_{2}^{3}{:}(G_{7}{:}G_{3}).$
@@
Let $\mathcal{M}$} be a map with square-free Euler characteristic, and let $G\leqslant{\sf Aut}{\mathcal{M}}$ be solvable and arc-transitive on $\mathcal{M}$}. Then, G=(A:B):KG=(A{:}B){:}K, where $A$ is abelian, $B$ is nilpotent, $A\cap{\bf Z}(A{:}B)=1$, and one of the following holds :
@@
    $\mathcal{M}$} is $G$-arc-transitive, $K=G_{2}$ is ${\sf Z}_{2}$ or one of the groups given in Proposition 4.2.
@@
    $\mathcal{M}$} is $G$-regular, $K=G_{\{2,3\}}$ is one of the groups :
    $X$, $X\times{\sf Z}_{2}$, ${\rm D}_{2.3^{\ell}}\times{\rm S}_{4}$, $({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})\times{\rm D}_{6}$,
    where $X\in\{{\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}},\,({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2},\,(({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})\times{\sf Z}_{3}){:}{\sf Z}_{2}\}.$
@@
    $\mathcal{M}$} is $G$-vertex-reversing, $K=G_{\{2,3\}}$ is either one of the groups given in (2), or one of the groups :
    $Y$, $Y\circ_{{\sf Z}_{2}}{\sf Z}_{4}$, $({\sf Z}_{3^{\ell}}{:}{\sf Z}_{4})\circ({\rm Q}_{8}{:}{\rm S}_{3})$, $({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})\circ({\sf Z}_{3}{:}{\sf Z}_{4})$,
    where $Y\in\{{\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}},\,({\sf Z}_{3^{\ell}}\times({\rm Q}_{8}{:}{\sf Z}_{3})){:}{\sf Z}_{2},\,({\sf Z}_{3}\times({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}\}.$
  • (4)
    $\mathcal{M}$} is $G$-vertex-rotary, and either $K=G_{\{2,3\}}$ or $G_{\{2,7\}}$ is one of the groups :
    ${\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}}$,  ${\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})$,  ${\sf Z}_{4}\circ_{{\sf Z}_{2}}({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})$,  ${\sf Z}_{2}^{3}{:}{\sf Z}_{7^{\ell}}$ ;
    or $K=G_{\{2,3\}}$ is one of the groups given in Theorem 1.1 (iii), excluding the following four :
    $({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2}$,  $({\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}$,  $({\sf Z}_{3^{\ell}}\times({\rm Q}_{8}{:}{\sf Z}_{3})){:}{\sf Z}_{2}$,  $({\sf Z}_{3}\times({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}$.
The proofs of the above two results are respectively given in Section 4 - 5 and Section 6. We remark that, Theorem 1.2 does not guarantee the existence of arc-transitive maps $\mathcal{M}$} corresponding to each of these groups. The existence and construction of such maps is the focus of the next paper in this series. Here, we construct several interesting examples of regular maps with square-free Euler characteristic in Section 3, from which the following result follows.
@@
There are infinite families of regular maps which have square-free Euler characteristic of form n(n2)-n(n-2) or n(n3)-n(n-3), with $n$ an positive integer.
@@ 2. Preliminaries
Let $G$ be a finite group, and let $H\leqslant G$ be a subgroup of $G$. Denote by $G^{\prime}$ the commutator subgroup of $G$, and by ${\bf N}_{G}(H)$ and ${\bf C}_{G}(H)$ the normalizer and the centralizer of $H$ in $G$, respectively. If $G$ is solvable, it is well-known that, for any subset $\pi$ of prime divisors of $|G|$, there exist Hall $\pi$-subgroups of $G$ with any two of them conjugate in $G$, and any $\pi$-subgroup is contained in some Hall $\pi$-subgroup.
@@
It is known (refer to [13]) that an arc-transitive (multi-) graph $\it\Gamma$} can be expressed as a coset graph :
${\it\Gamma}={\sf Cos}(G,H,J),\ \mbox{or}\ {\sf Cos}(G,H,HgH)\ \mbox{in the simple graph case},$
which has vertex set $[G:H]$ and edge set $[G:J]$, such that
    $Hg_{1}$ is incident with $Jg_{0}$ if and only if $g_{1}g_{0}^{-1}\in HJ$; or equivalently,
    $Hg_{1}$, $Hg_{2}$ are adjacent if and only if $g_{2}g_{1}^{-1}\in HgH$ in the simple graph case.
Further, $\it\Gamma$} is of valency $|H:H\cap H^{g}|$, where $g\in J\setminus(H\cap J)$, and $\it\Gamma$} is connected if and only if $G=\langle H,J\rangle$.
@@
A flag of $\mathcal{M}$} is an incidence triple $(\alpha,e,f)$ with $\alpha\in V$, $e\in E$ and $f\in F$. Let $G\leqslant{\sf Aut}({\mathcal{M}})$, and let $(\alpha,e,f)$ be a flag. It is well-known that, each of the stabilizers $G_{\alpha}$ and $G_{f}$ is cyclic or dihedral, and meanwhile, $G_{e}=1$, ${\sf Z}_{2}$ or ${\sf Z}_{2}^{2}$.
@@
(1) $\mathcal{M}$} is a $G$-regular map (type 1), which is determined by a regular triple $(x,y,z)$ for $G$ satisfying :
$\mbox{$G=\langle x,y,z\rangle$, $|x|=|y|=|z|=2$ and $\langle x,z\rangle={\sf Z}_{2}^{2}$}.$
In this case, $G={\sf Aut}({\mathcal{M}})$, the stabilizers $G_{\alpha}=\langle x,y\rangle,\,G_{e}=\langle x,z\rangle,\,G_{f}=\langle y,z\rangle,$ and the underlying graph $\it\Gamma$} is expressed as the coset graph ${\sf Cos}(G,H,HJ)$, where $H=\langle x,y\rangle$, $J=\langle x,z\rangle$, so $\it\Gamma$} is connected, of valency $|H:H\cap H^{z}|$.
@@
The following lemma established in [11] characterizes Sylow subgroups of ${\sf Aut}({\mathcal{M}})$ in the case where $\mathcal{M}$} is with square-free characteristic. Note that, if $\chi({\mathcal{M}})$ is square-free, then of course so is $\gcd(\chi({\mathcal{M}}),|{\sf Aut}{\mathcal{M}}|)$.
@@ 3. Examples of Square-free Euler Characteristic Regular Maps
We construct three infinite families of regular maps with square-free Euler characteristic in this section. The constructions are based on the following groups:
@@
The first family is given below.
@@
Let $X$ be the group given in ($\ast$), and let $G=\langle a,s\rangle\times\langle b,t\rangle={\rm D}_{2n}\times{\rm D}_{2n}\leqslant X$. Assume that $n$ is odd. Define $x,y,z$ as follows :
@@
Then, $(x,y,z)$ is a triple of involutions such that $xz=zx$. Note that,
$(yx)^{2}=(abt)^{2}=a^{2}$, $(yxz)^{2}=(abs)^{2}=b^{2}$,
and meanwhile, $\langle a\rangle=\langle a^{2}\rangle$, $\langle b\rangle=\langle b^{2}\rangle$ since $n$ is odd. Thus, $a,b,s$ and $t=xz$ are all contained in $\langle x,y,z\rangle,$ so $G=\langle x,y,z\rangle$, and $(x,y,z)$ is a regular triple for $G$.
@@
$H=\langle x,y\rangle=\langle s,abst\rangle=\langle abt\rangle{:}\langle s\rangle=\langle a,s\rangle\times\langle bt\rangle={\rm D}_{4n},$
and the face stabilizer $L$ is
@@
Thus, $(|V|,|E|,|F|)=(\frac{|G|}{|H|},\frac{|G|}{4},\frac{|G|}{|L|})=(n,n^{2},2n)$, and the Euler characteristic
$\chi({\mathcal{M}}_{n})=n-n^{2}+2n=-n(n-3)$.
Let $n=4x+1$. Then $\chi({\mathcal{M}}_{n})=-2(4x+1)(2x-1)$. By a result given by G. Ricci in 1933 (which is improved by $\rm P.Erd\ddot{o}s$} in [8]), there are infinitely many positive integers $x$ for which f(x)=(4x+1)(2x1)f(x)=(4x+1)(2x-1) is square-free, so there are infinitely many odd values of $n$ such that $\chi({\mathcal{M}}_{n})$ is square-free. Here are some examples :
$\begin{array}[]{|c|c|c|c|c|c|c|c|}\hline\cr n&5&13&17&29&37&41&\cdots\\ \hline\cr-n(n-3)&-2.5&-2.5.13&-2.7.17&-2.13.29&-2.17.37&-2.19.41$$\cdots\\ \hline\cr\end{array}$
At last, the valency of ${\mathcal{M}}_{n}$ is $\frac{2|E|}{|V|}=2n$, and the face length is $\frac{2|E|}{|F|}=n$. Note that, $H\cap H^{z}=\langle a,s\rangle\cong{\rm D}_{2n}$ is of index $2$ in $H$. Then, the underlying graph ${\it\Gamma}={\sf Cos}(G,H,HJ)$ is connected of valency 2, so it is the multiple cycle ${\bf C}_{n}^{(n)}$. $\Box$
The second family is given below.
@@
Let $X$ be the group given in ($\ast$), and let $G=\langle a,b\rangle{:}\langle st,\sigma\rangle={\sf Z}_{n}^{2}{:}{\sf Z}_{2}^{2}\leqslant X$. Define $x,y,z$ as follows :
@@
Then, $(x,y,z)$ is a triple of involutions such that $xz=zx$. Note that,
$zy=b$, $(zy)^{x}=b^{\sigma}=a$.
Thus, $a,b,\sigma$ and $st=a^{-1}y$ are all contained in $\langle x,y,z\rangle$, so $G=\langle x,y,z\rangle$, and $(x,y,z)$ is a regular triple for $G$.
@@
$H=\langle x,y\rangle=\langle ast\sigma\rangle{:}\langle\sigma\rangle={\rm D}_{4n},$
and the face stabilizer $L$ is
@@
Thus, $(|V|,|E|,|F|)=(\frac{|G|}{|H|},\frac{|G|}{4},\frac{|G|}{|L|})=(n,n^{2},2n)$, and the Euler characteristic
$\chi({\mathcal{M}}_{n})=n-n^{2}+2n=-n(n-3)$.
By Lemma 3.2, there are infinitely many odd values of $n$ such that $\chi({\mathcal{M}}_{n})$ is square-free. In a similar way, we can prove that there are infinitely many even values of $n$ with $n=4x+2$, such that $\chi({\mathcal{M}}_{n})=-2(2x+1)(4x-1)$ is square-free. Here are some examples :
$\begin{array}[]{|c|c|c|c|c|c|c|c|}\hline\cr n&2&10&14&22&26&34&\cdots\\ \hline\cr-n(n-3)&2&-2.5.7&-2.7.11&-2.11.19&-2.13.23&-2.17.31$$\cdots\\ \hline\cr\end{array}$
At last, the valency of ${\mathcal{M}}_{n}$ is $\frac{2|E|}{|V|}=2n$, and the face length is $\frac{2|E|}{|F|}=n$. Note that, $H\cap H^{z}=\langle a^{-1}b\rangle{:}\langle\sigma\rangle={\rm D}_{2n}$ is of index $2$ in $H$. Then, the underlying graph ${\it\Gamma}={\sf Cos}(G,H,HJ)$ is connected of valency 2, so it is the multiple cycle ${\bf C}_{n}^{(n)}$. $\Box$
@@
Let $G=X=\langle a,b\rangle{:}\langle s,\sigma\rangle={\sf Z}_{n}^{2}{:}{\rm D}_{8}$ be the group given in ($\ast$). Assume that $n$ is odd. Define $x,y,z$ as follows :
@@
Then, $(x,y,z)$ is a triple of involutions such that $xz=zx$. Note that,
$(zy)^{2}=(abt)^{2}=a^{2}$, $((zy)^{2})^{x}=(a^{2})^{x}=b^{2}$,
and meanwhile, $\langle a\rangle=\langle a^{2}\rangle$, $\langle b\rangle=\langle b^{2}\rangle$ since $n$ is odd. Thus, $a,b,s,\sigma$ are all contained in $\langle x,y,z\rangle$, so $G=\langle x,y,z\rangle$, and $(x,y,z)$ is a regular triple for $G$.
@@
$H=\langle x,y\rangle=\langle\sigma,s\rangle={\rm D}_{8},$
and the face stabilizer $L$ is